# Maximizing Relation Extraction Potential: A Data-Centric Study to Unveil Challenges and Opportunities

### v1

@@ Abstract
Relation extraction is a Natural Language Processing task aiming to extract relationships from textual data. It is a critical step for information extraction. Due to its wide-scale applicability, research in relation extraction has rapidly scaled to using highly advanced neural networks. Despite their computational superiority, modern relation extractors fail to handle complicated extraction scenarios. However, a comprehensive performance analysis of the state-of-the-art relation extractors that compile these challenges has been missing from the literature, and this paper aims to bridge this gap. The goal has been to investigate the possible data-centric characteristics that impede neural relation extraction. Based on extensive experiments conducted using 15 state-of-the-art relation extraction algorithms ranging from recurrent architectures to large language models and seven large-scale datasets, this research suggests that modern relation extractors are not robust to complex data and relation characteristics. It emphasizes pivotal issues, such as contextual ambiguity, correlating relations, long-tail data, and fine-grained relation distributions. In addition, it sets a marker for future directions to alleviate these issues, thereby proving to be a critical resource for novice and advanced researchers. Efficient handling of the challenges described can have significant implications for the field of information extraction, which is a critical part of popular systems such as search engines and chatbots. Data and relevant code can be found at https://github.com/anushkasw/MaxRE.
@@ I Introduction
Modern-day relation extractors heavily employ deep learning architectures such as BERT [6], ERNIE [7], and more recently, GPT-3 [8] for extracting complex entity and relation representations from the input data. Regardless, it has been observed that modern relation extractors perform well in simple scenarios but poorly on complicated data and relation types. For example, issues like longer input lengths, multiple relations per sample, overlapping entities, fine-grained relationships, and long-tail data distribution pose challenges that most relation extractors cannot handle efficiently. Although a plethora of traditional surveys that comprehensively summarize the literature exists in this field [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 18, 20, 21, 22, 23, 24, 25], they lack a thorough discussion of the issues faced by the relation extractors and their potential causes. Additionally, the perspective of finding drawbacks in the algorithms rather than concentrating on the data-specific reasons for the challenges is favored in the literature. More importantly, most works focus on very few datasets and algorithm combinations, making it difficult to get a holistic view of the attributes that pose complications for relation extractors [20, 21, 22, 23, 24, 25]. This also entails that the challenges and future directions proposed by such studies may lack credibility. Consequently, we perform an in-depth data-centric analysis to put the challenges in perspective and provide a navigable guide through the abundance of research in this field for novice readers to find problem areas that require more attention.
This paper presents the first data-centric performance analysis that investigates a diverse set of relation extraction algorithms to highlight the complex data attributes adversely affecting them. Compared to the existing studies discussed above, this work aims to investigate the root cause of the issues faced by the algorithms by focusing on the attributes of the data. The study summarizes the most important relation extraction paradigms in the literature and highlights the key challenges and future directions backed by an exhaustive set of experiments. Finally, the goal here is not to find shortcomings in the algorithms but rather to determine the data characteristics that make it difficult for neural relation extractors to detect relationships efficiently. This enables future research that addresses these challenges to help create better relation extractors. Thus, the contributions of this work are as follows:
@@
    Comprehensive algorithm comparison and sample level analysis of 15 neural relation extractors, including the much recent LLM-based algorithms on seven large-scale datasets to highlight reasons behind the performance gap in the field.
@@
    A software repository of datasets and re-implemented LLM-based algorithms shared for practitioners in industry and academia for reproducibility [^1] https://github.com/anushkasw/MaxRE.
@@ II-A1 Traditional Approaches
Extracting relationships between two entities requires an efficient understanding of relevant information that can be present anywhere in the input sample. Recurrent architectures, such as the LSTM, proved very useful due to their capability of capturing long-range dependencies. Thus, all initial relation classifiers employed LSTM architectures with a fully connected softmax layer for classification. Algorithms such as Attn-BLSTM [32] and CRNN-Att [33] were among the first to provide enhanced solutions to relation extraction by incorporating different strategies of attention mechanisms [34] enabling them to extract essential information present anywhere in the input sample. However, they did not account for the position and presence of the entities in the training process. Subsequent research focused on using entity-specific information, such as position features, to aid the classification process [35, 36]. However, the recurrent architectures lacked at relation classification due to their shorter context window making them prone to forgetting past states. Thus, with the introduction of the transformer architecture [34], pre-trained language models (PLMs) were heavily employed for relation classification.
Popular PLMs such as BERT and RoBERTa [37] have performed exceptionally well in various NLP tasks. They are trained on large corpora of data using masked language modeling (MLM) styled training and can be easily fine-tuned on smaller task-specific datasets, providing significant performance gains. Initially, PLM-based algorithms for relation classification aimed to understand contextual information with minimum knowledge of the entities and gradually progressed to more advanced forms of additional knowledge. For example, RBERT [38] and RoBERTa-large [39] used contextualized token representations from pre-trained BERT and RoBERTa models and incorporated knowledge about the global position of entities by introducing special tokens at the input level. Subsequently, algorithms started incorporating more advanced forms of knowledge by mapping entities to their knowledge base (KB) counterparts [7, 40, 41, 42]. The introduction of PLM-based algorithms significantly advanced the research in the field due to the capabilities of the networks to extract and connect contextual representations of the entities and the relations. However, such relation classifiers suffered due to the objective gap between the pre-training and fine-tuning strategies. Most PLM-based relation classifiers used MLM to pre-train the language model, while a classification objective was used for fine-tuning. Prompt-based algorithms were introduced to alleviate the reliance on separate classification modules and use the language models as a classifier.
@@ II-A2 Prompt-based Approaches
Prompt-based techniques aim to model the relation classification problem as an auto-regressive text generation task, enabling language models to be predictors rather than feature extractors [43]. This ability is achieved using cloze-style (fill-in-the-blanks) prompts or statements to probe the language model to perform various downstream tasks. A typical prompt consists of a template and verbalizing the label space. The prompt is coupled with the input sentence and made to answer a cloze-style statement. This technique is now widely employed in the NLP literature and has shown great improvements in zero and few-shot learning where training data is scarce. However, adapting the prompting strategy to relation classification has been a challenging task due to the diversity of the large label space. Not only is it difficult to create distinct verbalizations of multiple relations manually, but it is also difficult to fit the large label space in prompts, especially for expensive LLMs such as the GPT variants. Thus, prompt-based approaches for relation classification aimed to alleviate these issues by creating custom prompt-tuning strategies by employing PLMs like BERT, RoBERTa, T5, and BART [24, 44, 45].
@@ II-B Distant Supervision
The distant supervision paradigm was developed to cater to the need for larger annotated datasets by creating them without human intervention. This was accomplished by aligning training corpora with KBs such as Freebase [^4] https://en.wikipedia.org/wiki/Freebase_(database), DBPedia [^5] https://www.dbpedia.org/ and Wikidata [^6] https://www.wikidata.org/wiki/Wikidata:Main_Page. Mintz et al. [50] were the first to introduce this paradigm in the context of relation extraction. According to their research, if two entities (e1,e2) participate in a relation X present in a knowledge base D, then all sentences from a text corpus C containing (e1,e2) will participate in the relation X. This way, large amounts of data could be automatically annotated using relations from knowledge bases. Although this technique attracted much attention, it was built on a stringent assumption that all sentences containing (e1,e2) must belong to a given relation. This assumption was prone to falter, leading to the wrong labeling of the datasets.
Most research in distant supervision revolved around alleviating the problem of noisy labels by finding efficient ways of selecting sentences associated with specific relations. A prominent work done by Riedel et al. [51] relaxed the above assumption and restated the distant supervision paradigm as follows: If two entities (e1,e2) belonging to a text corpus C participate in a relation X present in a knowledge base D then out of all the sentences containing (e1,e2) from C, at least one sentence will participate in the relation X. This new assumption significantly helped reduce the number of noisy labels; however, it worked on the assumption that an entity pair could not participate in multiple relations. This assumption faltered when it came to large, distantly supervised datasets. Thus, the multi-instance multi-label (MIML) paradigm for relation extraction was introduced to enable the joint modeling of entities and relations and to combat the problem of overlapping relations[52, 53]. The MIML paradigm emphasized the need and benefits of extracting relations and entities in a single pipeline. Thus, the joint relation extraction paradigm was developed.
@@ II-C Joint Relation Extraction Algorithms
One of the primary challenges for the early joint relation extractors was the problem of multiple relations and overlapping entities. The assumption that two entities participate in a single relation in a given sentence is rarely satisfied in real-world data. Relationships can occur in various forms in natural text. For example, the sentence “Dominican Republic signer and songwriter Juan Luis Guerra led the Latin Grammy Awards nominations followed by Puerto Rico’s Ricky Martin” depicts the relation city of birth between the entities Dominican Republic and Juan Luis Guerra. However, the same relationship exists between Puerto Rico and Ricky Martin. A sentence can exhibit multiple relations by sharing the same or different entities [59]. Figure 3 shows a few more examples of this complex scenario. Entity pair overlap (EPO) and single entity overlap (SEO) are two instances of this problem where multiple relations are shared between both or a single entity. Thus, subsequent research in this field was done to develop relation extractors robust to this issue. Some methods to solve this problem included creating better representation schema for instances with overlap [60], incorporating relational knowledge as apriori to avoid extracting irrelevant entities [39], decomposition-based methods [61, 62] and using the seq2seq paradigm to circumvent the need for triplet extraction in a particular order [58, 58].
@@
This section highlighted the plethora of research endeavors in relation extraction literature. The efforts can be categorized into classification and extraction strategies that aim to classify relationships from a pre-defined label set or extract entity and relation triplets, respectively. The algorithms introduced have progressively advanced over time, each employing a significantly more advanced deep learning backbone than the previous one. The most recent one was the introduction of relation extraction algorithms using OpenAI’s GPT-3 architecture, which has shown notable performance in low-resource scenarios. These research efforts have imparted significant gains to the field of relation extraction. However, there is still a considerable performance gap in the field where relation extractors have not reached their full potential. This study hypothesizes that the disparity comes from input data characteristics that pose complicated use cases for relation extractors to tackle. Thus, the study highlights some of these complex data characteristics by looking at a diverse group of relation extraction algorithms and critically examining their effectiveness in extracting relationships from complex textual data. To this end, the next section discusses the methodology used for a data-centric performance analysis and some preliminary insights from it.
@@ III-A Datasets & Complex Data Characteristics
SemEval [^7] https://github.com/sahitya0000/Relation-Classification [65] 10718 19 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
NYT10888https://github.com/truthless11/HRL-RE/tree/master [51] 93598 29 $\checkmark$ $\checkmark$ $\checkmark$
FewRel [^9] https://thunlp.github.io/1/fewrel1.html [66] 27328 80 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CrossRE [^10] https://github.com/mainlp/CrossRE [67] 19761 17 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TACRED [^11] https://nlp.stanford.edu/projects/tacred/ [68] 106264 42 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
RETACRED [^12] https://github.com/gstoica27/Re-TACRED[69] 91467 40 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
WebNLG [^13] https://github.com/weizhepei/CasRel/tree/master [70] 14486 171 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ III-B1 Relation Classification Algorithms
The relation classification experiments were conducted using algorithms ranging from recurrent architectures to modern LLMs. Although LLMs provide significant performance gains for relation classification, it is essential to analyze recurrent algorithms to get a holistic view of the field. Furthermore, existing research only showcases the performance of these algorithms on a select few datasets. Hence, a comprehensive analysis of the recurrent algorithms for their generalizability and robustness to complex data characteristics might lead to novel and crucial insights into the field. Thus, this work utilizes three recurrent architectures.
@@
A crucial area with significant research has been the use of PLMs for relation classification. These algorithms aim to fine-tune SOTA PLMs by introducing various learning strategies. Some of the most high-performing algorithms that have been implemented for this work have been discussed below.
@@
  • ERNIE [7] incorporates information about entities by mapping them to their knowledge base (KB) counterparts. It uses a dual encoder strategy in which the first encoder (BERT) creates holistic input representations, and the second decoder injects external knowledge from the KBs.
@@
  • GPT-RE [74] utilizes sentence similarity (simCSE [75]) with a KNN-based demonstration retrieval strategy to perform in-context learning with a GPT-based architecture. The study also introduces the use of relation representations from PLMs, such as BERT, that are finetuned for the RE task to extract relevant demonstrations. However, due to the unavailability of the source code, only the sentence similarity paradigm was used in this study.

### v2

@@ Abstract
Relation extraction is a Natural Language Processing task that aims to extract relationships from textual data. It is a critical step for information extraction. Due to its wide-scale applicability, research in relation extraction has rapidly scaled to using highly advanced neural networks. Despite their computational superiority, modern relation extractors fail to handle complicated extraction scenarios. However, a comprehensive performance analysis of the state-of-the-art extractors that compile these challenges has been missing from the literature, and this paper aims to bridge this gap. The goal has been to investigate the possible data-centric characteristics that impede neural relation extraction. Based on extensive experiments conducted using 15 state-of-the-art relation extraction algorithms ranging from recurrent architectures to large language models and seven large-scale datasets, this research suggests that modern relation extractors are not robust to complex data and relation characteristics. It emphasizes pivotal issues, such as contextual ambiguity, correlating relations, long-tail data, and fine-grained relation distributions. In addition, it sets a marker for future directions to alleviate these issues, thereby proving to be a critical resource for novice and advanced researchers. Efficient handling of the challenges described can have significant implications for the field of information extraction, which is a critical part of popular systems such as search engines and chatbots. Data and relevant code can be found at https://aaig.ece.ufl.edu/projects/relation-extraction.
@@ I Introduction
Modern-day relation extractors heavily employ deep learning architectures such as BERT [6], ERNIE [7], and more recently, GPT-3 [8] for extracting complex entity and relation representations from the input data. Regardless, it has been observed that modern relation extractors perform well in simple scenarios but poorly on complicated data and relation types. For example, issues such as longer input lengths, multiple relations per sample, overlapping entities, fine-grained relationships, and long-tail data distribution pose challenges that most relation extractors cannot handle efficiently. Although a plethora of traditional surveys that comprehensively summarize the literature exists in this field [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 18, 20, 21, 22, 23, 24, 25], they lack a thorough discussion of the issues faced by the relation extractors and their potential causes. Additionally, the perspective of finding drawbacks in the algorithms rather than concentrating on the data-specific reasons for the challenges is favored in the literature. More importantly, most works focus on very few datasets and algorithm combinations, making it difficult to get a holistic view of the attributes that pose complications for relation extractors [20, 21, 22, 23, 24, 25]. This also entails that the challenges and future directions proposed by such studies may lack credibility. Consequently, we perform an in-depth data-centric analysis to put the challenges in perspective and provide a navigable guide through the abundance of research in this field for novice readers to find problem areas that require more attention.
This paper presents the first data-centric performance analysis that investigates a diverse set of relation extraction algorithms to highlight the complex data attributes that adversely affect them. Compared to the existing studies discussed above, this work aims to investigate the root cause of the issues faced by the algorithms by focusing on the attributes of the data. The study summarizes the most important relation extraction paradigms in the literature and highlights the key challenges and future directions supported by an exhaustive set of experiments. Finally, the goal here is not to find shortcomings in the algorithms but rather to determine the data characteristics that make it difficult for neural relation extractors to detect relationships efficiently. This enables future research that addresses these challenges to help create better relation extractors. Thus, the contributions of this work are as follows:
@@
    Comprehensive algorithm comparison and sample level analysis of 15 neural relation extractors, including the recent LLM-based algorithms on seven large-scale datasets to highlight reasons behind the performance gap in the field.
@@
    A software repository of datasets and re-implemented LLM-based algorithms shared for practitioners in industry and academia for reproducibility [^1] https://aaig.ece.ufl.edu/projects/relation-extraction.
@@ II-A1 Traditional Approaches
Extracting relationships between two entities requires an efficient understanding of relevant information that can be present anywhere in the input sample. Recurrent architectures, such as the LSTM, proved very useful due to their capability of capturing long-range dependencies. Thus, all initial relation classifiers employed LSTM architectures with a fully connected softmax layer for classification. Algorithms such as Attn-BLSTM [32] and CRNN-Att [33] were among the first to provide enhanced solutions to relation extraction by incorporating different strategies of attention mechanisms [34] enabling them to extract essential information present anywhere in the input sample. Yet, they did not account for the position and presence of the entities in the training process. Subsequent research focused on using entity-specific information, such as position features, to aid the classification process [35, 36]. However, the recurrent architectures struggled with relation classification due to their shorter context window making them prone to forgetting past states. Thus, with the introduction of the transformer architecture [34], pre-trained language models (PLMs) were heavily employed for this task.
Popular PLMs such as BERT and RoBERTa [37] have performed exceptionally well in various NLP tasks. They are trained on large corpora of data using masked language modeling (MLM) styled training and can be easily fine-tuned on smaller task-specific datasets, providing significant performance gains. Initially, PLM-based algorithms for relation classification aimed to understand contextual information with minimum knowledge of the entities and gradually progressed to more advanced forms of additional knowledge. For example, RBERT [38] and RoBERTa-large [39] used contextualized token representations from pre-trained BERT and RoBERTa models and incorporated knowledge about the global position of entities by introducing special tokens at the input level. Subsequently, algorithms started incorporating more advanced forms of knowledge by mapping entities to their KB counterparts [7, 40, 41, 42]. The introduction of PLM-based algorithms significantly advanced the research in the field due to the capabilities of the networks to extract and connect contextual representations of the entities and the relations. However, such relation classifiers suffered due to the objective gap between the pre-training and fine-tuning strategies. Most PLM-based relation classifiers used MLM to pre-train the language model, while a classification objective was used for fine-tuning. Prompt-based algorithms were introduced to alleviate the reliance on separate classification modules and use the language models as a classifier.
@@ II-A2 Prompt-based Approaches
Prompt-based techniques aim to model the relation classification problem as an auto-regressive text generation task, enabling language models to be predictors rather than feature extractors [43]. This ability is achieved using cloze-style (fill-in-the-blanks) prompts or statements to probe the language model to perform various downstream tasks. A typical prompt consists of a template and verbalizations of the label space. The prompt is coupled with the input sentence and made to answer a cloze-style statement. This technique is now widely employed in the NLP literature and has shown great improvements in zero and few-shot learning where training data is scarce. However, adapting the prompting strategy to relation classification has been a challenging task due to the diversity of the large label space. Not only is it difficult to create distinct verbalizations of multiple relations manually, but it is also difficult to fit the large label space in prompts, especially for expensive LLMs such as the GPT variants. Thus, prompt-based approaches for relation classification aimed to alleviate these issues by creating custom prompt-tuning strategies by employing PLMs like BERT, RoBERTa, T5, and BART [24, 44, 45].
@@ II-B Distant Supervision
The distant supervision paradigm was developed to cater to the need for larger annotated datasets by creating them without human intervention. This was accomplished by aligning training corpora with KBs such as Freebase [^4] https://en.wikipedia.org/wiki/Freebase_(database), DBPedia [^5] https://www.dbpedia.org/ and Wikidata [^6] https://www.wikidata.org/wiki/Wikidata:Main_Page. Mintz et al. [50] were the first to introduce this paradigm in the context of relation extraction. According to their research, if two entities (e1,e2) participate in a relation X present in a KB D, then all sentences from a text corpus C containing (e1,e2) will participate in the relation X. This way, large amounts of data could be automatically annotated using relations from KBs. Although this technique attracted much attention, it was built on a stringent assumption that all sentences containing (e1,e2) must belong to a given relation. This assumption was prone to falter, leading to the wrong labeling of the datasets.
Most research in distant supervision revolved around alleviating the problem of noisy labels by finding efficient ways of selecting sentences associated with specific relations. A prominent work done by Riedel et al. [51] relaxed the above assumption and restated the distant supervision paradigm as follows: If two entities (e1,e2) belonging to a text corpus C participate in a relation X present in a KB D then out of all the sentences containing (e1,e2) from C, at least one sentence will participate in the relation X. This new assumption significantly helped reduce the number of noisy labels; however, it worked on the assumption that an entity pair could not participate in multiple relations. This assumption faltered when it came to large, distantly supervised datasets. Thus, the multi-instance multi-label (MIML) paradigm for relation extraction was introduced to enable the joint modeling of entities and relations and to combat the problem of overlapping relations[52, 53]. The MIML paradigm emphasized the need and benefits of extracting relations and entities in a single pipeline. Consequently, the joint relation extraction paradigm was developed.
@@ II-C Joint Relation Extraction Algorithms
One of the primary challenges for the early joint relation extractors was the problem of multiple relations and overlapping entities. The assumption that two entities participate in a single relation in a given sentence is rarely satisfied in real-world data. Relationships can occur in various forms in natural text. For example, the sentence “Dominican Republic signer and songwriter Juan Luis Guerra led the Latin Grammy Awards nominations followed by Puerto Rico’s Ricky Martin” depicts the relation city of birth between the entities Dominican Republic and Juan Luis Guerra. However, the same relationship exists between Puerto Rico and Ricky Martin. A sentence can exhibit multiple relations by sharing the same or different entities [59]. Figure 3 shows a few more examples of this complex scenario. Entity pair overlap (EPO) and single entity overlap (SEO) are two instances of this problem where multiple relations are shared between both or a single entity. Thus, subsequent research in this field was done to develop relation extractors robust to this issue. Some methods to solve this problem included creating better representation schema for instances with overlap [60], incorporating relational knowledge as apriori to avoid extracting irrelevant entities [39], decomposition-based methods [61, 62] and using the seq2seq paradigm to circumvent the need for triplet extraction in a particular order [58].
@@
This section highlighted the plethora of research endeavors in relation extraction literature. The efforts can be categorized into classification and extraction strategies that aim to classify relationships from a pre-defined label set or extract entity and relation triplets, respectively. The algorithms introduced have progressively advanced over time, each employing a significantly more advanced deep learning backbone than the previous one. The most recent one was the introduction of relation extraction algorithms using OpenAI’s GPT-3 architecture, which has shown notable performance in low-resource scenarios. These research efforts have imparted significant gains to the field of relation extraction. However, there is still a considerable performance gap in the field where relation extractors have not reached their full potential. This study hypothesizes that the disparity comes from input data characteristics that pose complicated use cases for relation extractors to tackle. Thus, the study highlights some of these complex data characteristics by looking at a diverse group of relation extraction algorithms and critically examining their effectiveness in extracting relationships from complex textual data. To this end, the next section discusses the methodology used for the data-centric performance analysis and some preliminary insights from it.
@@ III-A Datasets & Complex Data Characteristics
SemEval [65] 10718 19 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
NYT10 [51] 93598 29 $\checkmark$ $\checkmark$ $\checkmark$
FewRel [66] 27328 80 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CrossRE [67] 19761 17 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TACRED [35] 106264 42 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
RETACRED [68] 91467 40 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
WebNLG [69] 14486 171 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ III-B1 Relation Classification Algorithms
The relation classification experiments were conducted using algorithms ranging from recurrent architectures to modern LLMs. Although LLMs provide significant performance gains for relation classification, it is essential to analyze recurrent algorithms to get a holistic view of the field. Furthermore, existing research only showcases the performance of these algorithms on a select few datasets. Hence, a comprehensive analysis of the recurrent algorithms for their generalizability and robustness to complex data characteristics might lead to novel and crucial insights into the field. Thus, this work utilizes three recurrent algorithms.
@@
A crucial area with significant research has been the use of PLMs for relation classification. These algorithms aim to fine-tune SOTA PLMs by introducing various learning strategies. The high-performing algorithms that have been implemented for this work are discussed below.
@@
  • ERNIE [7] incorporates information about entities by mapping them to their KB counterparts. It uses a dual encoder strategy in which the first encoder (BERT) creates holistic input representations, and the second decoder injects external knowledge from the KBs.
@@
  • GPT-RE [73] utilizes sentence similarity (simCSE [74]) with a KNN-based demonstration retrieval strategy to perform in-context learning with a GPT-based architecture. The study also introduces the use of relation representations from PLMs, such as BERT, that are finetuned for the relation extraction task to extract relevant demonstrations. However, due to the unavailability of the source code, only the sentence similarity paradigm was used in this study.